rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_red
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_red
icon_blue
icon_red
icon_orange
icon_red
icon_red
icon_red
icon_red
icon_red
icon_red
icon_red
icon_green
icon_red
icon_green
icon_green
 

FS#6205 — FS#10110 — LNS-1-PAR

Attached to Project— ADSL / SDSL
Incident
Backend / Core
CLOSED
100%
SFR collect on lns-1-par is down, we are investigating it with the manufacturer.
The customers were automatically switched onto the backup LNS.
Date:  Thursday, 06 February 2014, 16:02PM
Reason for closing:  Done
Comment by OVH - Thursday, 30 January 2014, 10:34AM

Following the return of the SFR collect to the Paris LNS during the night, instabilities have been detected.
The port was shutdown, access is on the backup LNS again.


Comment by OVH - Thursday, 30 January 2014, 11:44AM

We have changed the main LNS configuration LNS for lns-1-par customers to lns-1-lyo with lns-1-rbx as a backup, so that the sessions don't come back onto the Paris LNS, which is causing the instabilities.
The PPP sessions will thus be the Lyon LNS at the time of the next reset.


Comment by OVH - Friday, 31 January 2014, 11:40AM

We carried out tests with SFR without managing to find the anomaly, we are now working with Ericsson to find the source of the problem.


Comment by OVH - Friday, 31 January 2014, 17:58PM

We found a bug on the Ericsson router, we will receive the patch next week.
We will migrate the remaining FT customers onto the Roubaix LNS, and we will upgrade the node when we get the patch.


Comment by OVH - Thursday, 06 February 2014, 15:59PM

We have received the necessary patch and we will update the LNS tonight at around 0:00
http://status.ovh.co.uk/?do=details&id=6263